(x2 + 34)°, (6x + 4)°, and (10x + 10)°. Which name best describes this triangle?
Answer by stanbon(75887)

(x^2 + 34)°, (6x + 4)°, and (10x + 10)°.
Which name best describes this triangle?
------
Equation:
x^2+34 + 6x+4 + 10x+10 = 180
x^2 + 16x + 48 = 180
=====
x^2 + 16x - 132 = 0
----
(x-6)(x+22) = 0
-----
Positive solution:
x = 6
-----
x^2 + 34 = 36+34 = 70 degrees
6x+4 = 40 degrees
10x+10 = 70 degrees
=========================
Ans: isosceles triangle
==================
Cheers,
Stan H.